OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

jimthompson5910 (jim_thompson5910):

Solve the first equation for y 2x + y = 1 y = 1 - 2x y = -2x + 1

jimthompson5910 (jim_thompson5910):

Then plug this into the second equation and solve for x 4x + 2y = −2 4x + 2(-2x + 1) = −2 4x - 4x + 2 = -2 I'll let you finish

OpenStudy (anonymous):

But Beyonce had the best video of all time.

OpenStudy (anonymous):

so its x = 2 @jim_thompson5910

OpenStudy (anonymous):

what? O.o @Eleven17

jimthompson5910 (jim_thompson5910):

yeah i have no idea either...

jimthompson5910 (jim_thompson5910):

4x - 4x + 2 = -2 0x + 2 = -2 0+2 = -2 2 = -2 is that true?

OpenStudy (anonymous):

im not sure but i think so @jim_thompson5910

jimthompson5910 (jim_thompson5910):

2 is NOT equal to -2

jimthompson5910 (jim_thompson5910):

so 2 = -2 is NOT true

jimthompson5910 (jim_thompson5910):

therefore, there are NO solutions

jimthompson5910 (jim_thompson5910):

there's no set of numbers for x and y that will solve the original system

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!!!!!!!!!! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome